Show your solution using the critical value method.


a previous study on the effect of blended learning module on senior high school student's achievement revealed that mean score pi is 85 with standard deviation of 3. juliana, a senior high school statistic teacher, surveyed a class of 30 students test the hypothesis at a 0.10 level of significance


1
Expert's answer
2022-05-18T09:48:51-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=85"

"H_a:\mu\not=85"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=29" degrees of freedom, and the critical value for a two-tailed test is "t_c =1.699127."

The rejection region for this two-tailed test is "R = \{t: |t|>1.699127\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{83-85}{4\sqrt{30}}\approx-2.7386"

Since it is observed that "|t|= 2.7386 >1.699127= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed "df=29" degrees of freedom, "t=-2.7386" is "p =0.00617," and since "p=0.00617<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 85, at the "\alpha = 0.10" significance level.